Using Charpit’s method, find a complete integral of the following equations: xpq + yq^2 = 1


1
Expert's answer
2021-01-19T04:26:36-0500

f(x,y,z,p,q)=xpq+yq2-1=0,

"\frac{dp}{\frac{df}{dx}+p\frac{df}{dz}}=\frac{dq}{\frac{df}{dy}+q\frac{df}{dz}}=\frac{dz}{-p\frac{df}{dp}-q\frac{df}{dq}}=\frac{dx}{-\frac{df}{dp}}=\frac{dy}{-\frac{df}{dq}}"

"\frac{dp}{pq}=\frac{dq}{q^2}=\frac{dz}{-pxq-2yq^2-qxp}=\frac{dx}{-xq}=\frac{dy}{-xp-2yq}"

"\\frac{dq}{q}+ \\frac{dx}{x}=0 ,\nq\ndq\n\u200b\t\n + \nx\ndx\n\u200b\t\n =0"

ln(q)+ln(x)=a

ln(qx)=a

q=ea/x

"\frac{dp}{p}- \frac{dq}{q}=0"

ln(p)-ln(q)=b

ln(p/q)=b

p=ebq=e(a+b)/x

dz=pdx+qdy

dz=(e(a+b)/x)dx+(ea/x)dy",\int dz=\int e^{a+b}/x dx+\int e^a/x dy"


z=e(a+b)ln(x)+(eay/x)+c